# ToDo (#1): #2

### v2

@@ 3.4 Vacuum Chamber
The vacuum enclosure is composed of 5 parts : the upper and lower internal disk and 3 vacuum chambers. The vacuum sealing is done using 2 metallic O-rings per interface (from Omniseal [https://www.omniseal-solutions.com/] for example) except for the radial interface where Polymer O-rings could be used. The interface for the accessories will be done using a sleeve design and radial O-ring, described in the figure hereunder.
@@ 3.5 Coil configuration
[…]
@@
[…]
@@ 3.6 Cyclotron RF System
Table 3.4 provides a detailed description of the RF system. The system for IsoDAR cyclotron is divided into subsystems as follows:
  • The RF cavities allowing a fast acceleration to the final particle energy and keeping a large inter-turn separation. The cavities are triangular shaped and are located in the magnet valleys. There are 4 accelerating electrodes (Dees).
  • The RF tuning system that will compensate for thermal drift of cavities and maintain a stable resonating frequency.
  • The RF amplifiers that provide the necessary power to create the required RF electric field inside the cavity and the power to accelerate the beam.
@@
Dee voltage in the central region $70$ kV
@@
Dee angular extension ${}^{\circ}$
Cavity height 1800 mm
@@
Power dissipated per cavity 116 kW
@@
We consider these in the Sections below.
@@ 3.7 RF Cavities (Dees, Stems and Liners)
Figure 3.15: Early RF Dee computation – color scheme represent current density
The RF cavities of the cyclotron have been calculated using CST microwave studio [43] in order to define the shape of the Dee and the Dee stems. They have then been optimized also with CST from this early design. The optimization work is intended to reach the right resonance frequency, the Dee voltage Law, together with the minimal power dissipation and hence the highest quality factor (Q-factor). The field maps allow also the calculation of the couplers and voltage pick-up but this last calculation is yet to be performed.
The Dees and cavity walls are all made of OFHC copper that combines good electrical conductivity and good solder ability. The Dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm. The four round stems support the Dee and provide the right inductance in order to get the required resonance frequency. Their size and location allow reaching the required Dee voltage profile versus radius. Each Dee will be equipped with vacuum pumping panels to efficiently pump the median plane. One Dee is equipped with one electrostatic deflector for extraction. This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the Dee.
@@ 3.7.1 Integration of preliminary central region in RF Cavities
In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
  • The central region 3D model was limited to 200 mm from centre.
  • Both the Dee and counter dee original central region was removed by trimming the cavities at 800 mm from centre.
@@
These lofted transition leads to an acceptable linear vertical gap opening from 26 mm at a radius of 200 mm, to 50 mm at a radius of 800 mm. In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil. Symmetry with regards median plane symmetry was maintained. Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
@@ 3.7.2 RFQ Integration in RF Cavities
Some clearance provision has been set for the integration of the RFQ, ass seen on figure 3.19
@@ 3.7.3 RFQ getter pumps in RF Cavities
Mastering the quality of the high vacuum is key to the success of high-power beam cyclotrons. Incorporating getter pumps in the dees, in the case of IsoDAR, would allow to enhance vacuum quality. By integrating three housing for SAES UHV1400 getter pumps per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system. The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration. Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity. The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
@@ 3.7.4.1 Performance requirement target
  • Reference accelerating voltages: 70 kV @ 110 mm, 230 kV @ 2000 mm
@@ 3.7.4.2 Parametric optimization of the design
CST studio suite offers an efficient optimization tools set. After some initial exploration of the values of the above-mentioned design parameters, we have optimized them using the CST optimizer.
@@
Ratio between voltage at 110 mm and 2000mm 0.304348 0.303991
Mode1 frequency divided by 32.8MHz 1 0.999871
@@
Parameter Value [mm] Description
Deflector offset (D) 5.67 Offset of upper and lower face of deflector
Increase_diameter (B) -3.81 Change of diameter of the pillar
Move_pillar (A) -124.1 Radial displacement of the pillar
Back_face_offset (C) 4 Outward move of the back of the deflector
@@ 3.7.5 RF cavity computation results
  • Frequency: The RF cavity operates at a frequency of 32.796 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.303991, indicating that computed voltages at center would be 69.9 kV for a chosen voltage of 230 kV at the outer radius (2m),
  • Quality Factor (Q): 9617,
  • Stored Energy (for max RF of 230 kV at 2 m): 3.8 Joules
  • CST Computed Losses: The computed losses per cavity using CST simulations are 81.3 kW,
  • Estimated Power Dissipation: By experience, senior RF designers at IBA state that the CST eigenmode simulations underestimate the actual losses by up to 30%. If we increase the CST computed losses (81.3 kW) by this percentage, the actual estimated power dissipation per cavity becomes 116.2 kW.
  • Estimated Power Requirement: To effectively drive the Cyclotron, each cavity will require an estimated power input of at least 266.2 kW, which include power dissipation and beam loading.
  • Maximal Surface Current: The maximal surface current reaches 24324.6 A/m in some reduced locations. Which leads to a maximum surface power of 88.4 W/cm². This maximal surface power is localized on the dee edges. Careful attention will be required for thermal management in these regions. For the rest of the surfaces, the thermal dissipation is well below 10 W/cm². See figure 3.26 and 3.27 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.22, 3.23, 3.24 and 3.25.
@@ 3.7.5.1 Detailed power analysis
[…]
The ratio of the desired voltage to the reference voltage at 2m is 1.947. We can scale the currents and voltages using this factor. We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3). We thus have:
@@
Table 3.7: Optimized cavities, functional characteristics
@@
Stored energy [J] 1.001 3.796
Voltage at on the 2m orbit kV] 118.12 230
Voltage at center [kV] 35.91 69.9
Maximal voltage at outer edge of the dee [kV] 126.55 246.4
Computed losses [kW] 21.452 81.3
Maximal surface current [A/m] 12492.1 24324.6
@@
Angular Frequency $\omega$ $2\pi\times 32.796\times 10^{6}$ rad/s
Skin Depth $\delta$ $11.53$ µm
@@ 3.7.5.2 Radial voltage dependency
center 69.9
110 69.918
200 68.04
300 69.149
400 69.187
500 73.054
600 78.677
700 84.162
800 91.511
900 102.353
1000 111.984
1100 123.881
1200 138.013
1300 151.364
1400 162.731
1500 177.622
1600 189.519
1700 203.636
1800 212.224
1900 222.408
2000 230
Outer radius 246.4
@@ 3.7.6 RF Tuning System
The tuning system is a movable plate or loop that corrects the resonance frequency drift mainly caused by RF heating. The system here is based on four separated tuning plates (one per cavity) and located in the median plane. The plate is accurately displaced by an electrical motor via a ball screw. Similar systems are already used many commercial cyclotrons.
@@ 3.7.7 RF Amplifiers
The traditional way to produce such a high RF power is still today the use of vacuum tubes (triodes or tetrodes) and this is especially true in the accelerator world where the load constraints are more severe than simple antennas. On the other hand solid state technology made a significant step forward in the past years in terms of power density and robustness. Even though the last generation of LDMOS allows to make solid state amplifiers at nearly competitive price with respect to vacuum tubes and that are also very reliable, we do believe vacuum tubes still present reliability and maintainability advantages for powering the isoDAR cyclotron, given the harsh environment where those amplifiers will be located. Tubes amplifiers will also accommodate easily to the rapid variations of load due to beam temporal pattern.
@@ 3.7.8 RF Lines and Couplers
The choice between these two alternatives is mainly driven by the available space, the field patterns and the cooling concerns. The usual experience in this field is that both systems can work and couplers up to 800 kW at 107 MHz have been successfully built and operated.
Four couplers are needed for the complete cyclotron. The 300 kW power will be injected via one coupler per cavity.
Eventhough capacitive coupling might be less prone to multipacting by the fact that the coupling antenna could be easily DC biased, our expertise in high power coupler design advocates for an inductive solution. The design of inductive coupler being easier to watercool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
We have initiated a preliminary design computation of such a inductance coupler, feeding the RF power from a EIA 6”1/8 port. The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling. See figure 3.29. The adaptation tuning will be performed by rotation of the loop around the vertical axis. The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 362 mm, wich provides a good matching (figure 3.30) and a $S_{1,1}$ scattering parameter inferior to 0.01.
@@ 3.7.9 LLRF Electronics
The LLRF (Low Level RF) controls the RF amplitude and phase of the signal sent to the amplifiers. It also controls the tuning mechanism in order to maintain constant resonant frequencies in the cavities. The amplitude and phase control is based on feed-back signals coming from the cavities and amplifiers. In order to avoid the multipactor effect the LLRF manages also a pulsing mode that provides for an easy start-up and reduces the power dissipated in the amplifiers. One LLRF system per cavity is planned.
@@
3.8 Cyclotron magnetic performance
In this section we describe the latest modifications applied to the cyclotron regarding its magnetic properties. We will compare the results with the model from the Conceptual Design Report (CDR) from [6] to ensure its magnetic performance remain within specifications.
With the latest design of the vacuum chamber, see section 3.4, the pole radius was slightly reduced. To keep the magnetic field isochronous up to extraction at 120 MeV, the outer radius was therefore somewhat modified, see left image in Fig. 3.31. First, the curvature was modified to follow the curvature of the last orbit. This enhances the total flux in the region where needed. Second, the top layer of the pole at its outer radius will be made of permendur. This alloy has better magnetic properties at high field, and enhances again the magnetic flux. Finally, the pole height is increased by a slanted ridge at its extremity. This reduction of the pole gap should have no effect on the transmission and can be tuned to improve the dB/dr for the last orbits.
The steel used for the cyclotron yoke and poles is the SNVA6342, also used for example in the C400 cyclotron (see Ref. [44]). Its BH-curve is shown in fig 3.31. Here, we also see the better performance of the permendur alloy.
3.8.1 Isochronization

### v3

@@ 3.5 Vacuum Chamber
The vacuum enclosure comprises five parts: the upper and lower internal disk and 3 vacuum chambers. Fig 3.22 shows the vacuum enclosure in detail and the dimensions are in Table 3.5. The vacuum sealing is done using 2 metallic O-rings per interface (e.g., from Omniseal [PrecisionPolymerMetal]) except for the radial interface where Polymer O-rings could be used. The interface for the accessories will be done using a sleeve design and radial O-ring, described in Fig. 3.22.
@@
3.5.1 Risks and Mitigation
Risk: Achieving the required vacuum level in the presence of the beam. The calculated required vacuum is $5\cdot 10^{-8}$ mbar. Before the beam is turned on, the pumping plan provided will reach this with little difficulty, assuming there are no vacuum leaks. However, introduction of beam, with gas-desorption from surfaces hit by stray beam particles, could raise the pressure above the acceptable level.
Mitigation: Beam losses from other sources will be minimized. In addition, options for further surface cleaning and treatment, as well as the possible addition of more pumping capacity, should be developed.
@@ 3.6 Coil configuration
The CDR anticipated a slightly shorter coil than we now anticipate using. Specifically, the coil height has been increased from 200 to 240 mm. Its radial width remains at 250 mm. The total current used in the calculations is 180,850 amp-turns in each of the half coils. The apparent current density is thus 3.01 A/mm2, i.e. $5\%$ lower than the 3.17 A/mm2 foreseen in the CDR. For now, no extraction channels are anticipated in the return yoke. If they are added, an increase in main coil current will be required
@@
In Fig. 3.23 a schematic view of the coil is given. Also, a possible configuration of winding and cooling is calculated. Although this configuration does not exactly sum to a cross-section of 240 by 250 mm2, it shows that a cooling configuration is feasible.
@@ 3.7 Cyclotron RF System
The RF system for the IsoDAR cyclotron produces the primary acceleration of the ion beam and can be divided into the following subsystems:
  • The RF cavities which accelerate the beam and provide large inter-turn separation. The cavities are wedge-shaped and are located in the magnet valleys. There are 4 such cavities. The inner part is often called a dee whereas the outer part is called a dummy-dee.
  • The RF tuning system that will compensate for the thermal drift of the cavities and maintain a stable resonating frequency.
  • The RF amplifiers that provide the necessary power to create the required RF electric field inside the cavity.
@@
Dee voltage in the central region $65$ kV
@@
Acceleration gap angle 42
Cavity height 1812 mm
@@
Power dissipated per cavity 113 kW
@@
We consider these in the Sections below. The numerical parameters of the RF system are listed in Table 3.6.
@@ 3.8 RF Cavities (Dees, Stems and Liners)
The RF cavities of the cyclotron have been designed using CST microwave studio [cst:microwave_studio] in order to define the shape of the Dee and the Dee stems. We optimized our designs to reach the right resonance frequency, satisfy the Dee voltage Law, and minimize power dissipation. Ultimately, this improves the quality factor (Q-factor).
Figure 3.24: Section view of the RF cavity.
Figure 3.25: Outer view of the RF cavity.
The Dees and cavity walls are all made of oxygen-free high-conductivity (OFHC) copper, which has good electrical conductivity and good solderability. The Dees are fabricated out of a 20 mm thick copper plate and are made rigid by a supporting arm. The four round stems support the Dee and provide the right inductance in order to produce the required resonance frequency. Their size and location produce the desired Dee voltage vs. radius profile. Each Dee will be equipped with vacuum pumping panels to efficiently pump the median plane. One Dee is equipped with one electrostatic deflector for extraction. This fact has not been taken into account in detail in the RF simulations to date, but some room for housing the deflector has been provisioned at the back of the Dee. Figures 3.24 and 3.25 present the main elements composing the RF cavities.
@@ 3.8.1 Integration of preliminary central region in RF Cavities
In order to compute a more accurate voltage radial dependency law, we have integrated the central region design proposed by AIMA [aima:central] into the RF cavity modelling, several technical choices were made to ensure the effective implementation of this innovative configuration:
  • The central region design was trimmed at 150 mm from centre.
  • Both the Dee and counter dee original central region was removed by trimming the cavities at 250 mm from centre.
@@
The figure 3.26 shows the final integration of the central region. The lofted transitions can be better seen from top at figure 3.27. They leads to an acceptable linear vertical gap opening from 26 mm at a radius of 150 mm, to 50 mm at a radius of 250 mm. The section view of the model, presented at figure 3.28, shows the radial evolution of the vertical gap. In addition, adequate clearance provisions were incorporated, specifically to fit the inflector and optics coil. Symmetry with regard to the median plane symmetry was maintained. Power bridges between dees in the central regions were removed since we recommend to power the 4 cavities independently.
@@ 3.8.2 RFQ Integration in RF Cavities
Some clearance provision has been set for the integration of the RFQ, as seen in figure 3.29
@@ 3.8.3 RFQ getter pumps in RF Cavities
Incorporating getter pumps into the Dees allows us to improve our vacuum quality. By integrating three housing for SAES UHV1400 getter pumps [ModulesUHV] per dee, we aim to ensure highly efficient and reliable vacuum performance within the Cyclotron system. The use of six of those getter pumps per cavity facilitates the effective removal of residual gases, since achieving and maintaining ultra-high vacuum conditions essential for optimal particle acceleration. The figure 3.30 shows acceptable locations for three getter pump housings. Furthermore, our preliminary assessments suggest that it may be feasible to scale up the configuration to include up to two sets of five getter pumps per cavity. The very low electrical field of the locations of the housings will allow the getter pumps to function safely.
@@ 3.8.4.1 Performance requirement target
  • Reference accelerating voltages: 65 kV at 80 mm (capture), 230 kV at 2000 mm (extraction)
@@ 3.8.4.2 Parametric optimization of the design
We did some initial exploration of the parameter space, but we did the final optimization using the built-in CST optimizer.
@@
Ratio between voltage at 80 mm and 2000mm 0.28 0.283
Mode1 frequency [MHz] 32.8 32.805
@@
Design parameter Value [mm]
First pillar radialposition (A) 444.6
First pillar diameter (B) 117.7
Distance between dee and cavity back wall (C) 42
Total cavity height (D) 1412
@@ 3.8.5 RF cavity computation results
  • Frequency: The RF cavity operates at a frequency of 32.805 MHz, ensuring compatibility with the designed acceleration, and the RFQ integration,
  • Voltage Ratio (center to outer radius): The reached voltage ratio is 0.283, indicating that computed voltages at center would be 65.2 kV for a chosen voltage of 230 kV at the outer radius (2m),
  • CST computed quality factor (Q): 10586,
  • Estimated quality factor (Q): 7281,
  • Stored Energy (for max RF of 230 kV at 2 m): 4.06 Joules
  • CST Computed Losses: The computed losses per cavity using CST simulations are 79 kW,
  • Estimated Power Dissipation: By experience, senior RF designers at IBA state that the CST eigenmode simulations underestimate the actual losses by up to 30%. If we increase the CST computed losses (79 kW) by this percentage, the actual estimated power dissipation per cavity becomes 113 kW.
  • Estimated Power Requirement: To effectively drive the cyclotron, each cavity will require an estimated power input of at least 263 kW, which include power dissipation and beam loading.
  • Maximal Surface Current: The rms surface current reaches a maximum of 17052.6 A/m in some reduced locations. Which leads to a maximum surface power of 43.5 W/cm². This maximal surface power is localized on the dee edges. Careful attention will be required for thermal management in these regions. For the rest of the surfaces, the thermal dissipation is well below 10 W/cm². See figure 3.36 and 3.37 While matching the requirement, these computation results provide valuable information for next steps of the Cyclotron’s RF cavity design including consideration in managing thermal effects.
Some illustrations of the electrical and magnetic fields computed by CST are given for reference at figures 3.32, 3.33, 3.34 and 3.35.
@@ 3.8.5.1 Detailed power analysis
Figure 3.37: Surface current generating more than 10 W/cm2
The ratio of the desired voltage to the reference voltage at 2m is 2.013. We can scale the currents and voltages using this factor. We can also scale the energies, power and losses using this factor squared (formulae 3.1, 3.2 and 3.3). We thus have:
@@
Table 3.9: Optimized cavities, functional characteristics.
@@
Stored energy [J] 1 4.06
Voltage at on the 2m orbit kV] 114.2 230
Voltage at center [kV] 32.4 65.2
Maximal voltage at outer edge of the dee [kV] 124.6 250.9
Computed losses [kW] 19.48 78.9
Maximal surface current [A/m] 8470 17052.6
@@
Angular Frequency $\omega$ $2\pi\times 32.805\times 10^{6}$ rad/s
Skin Depth $\delta$ $11.5$ µm
@@ 3.8.5.2 Radial voltage dependency
Capture (80) 65.17
100 64.07
200 63.95
300 64.60
400 63.76
500 66.37
600 71.30
700 77.60
800 85.50
900 95.67
1000 107.09
1100 119.72
1200 133.04
1300 147.22
1400 160.77
1500 174.31
1600 187.84
1700 199.73
1800 211.94
1900 222.49
Extraction (2000) 230.00
back of dee 250.91
@@ 3.8.6 RF Tuning System
The tuning system is a movable plate that corrects the resonance frequency drift mainly caused by RF heating. The system here is based on four separated tuning plates (one per cavity) and located in the median plane. These plates are located at the back of the Dee and are controlled via a ball screw that is driven by a motor. This system is similar to that of many commercial cyclotrons.
@@ 3.8.7 RF Amplifiers
We plan to use vacuum tubes (triodes or tetrodes) to produce the RF power. We have considered using LDMOS amplifiers given the recent improvements in power density, reliability, and cost. However, we believe that vacuum tubes will perform better in the harsh environment and will be able to better accommodate the rapid variations in load due to the beam temporal patterns.
@@ 3.8.8 RF Lines and Couplers
Both of these systems are capable of providing the necessary 300 kW power to each cavity. Capacitive coupling is less prone to multipacting because the coupling antenna could be easily DC biased. But, our expertise in high power coupler design advocates for an inductive solution. The design of inductive coupler is easier to water-cool, and there is a wide area where H-field is high enough for efficient coupling behind the second pillar of the cavity.
We have initiated a preliminary design computation of such an inductance coupler, feeding the RF power from a EIA 6”1/8 port. The suggested design is a loop made out of a 20x60 mm rectangular section bar, offering room for water-cooling. See figure 3.39. The adaptation tuning will be performed by rotation of the loop around the vertical axis. The design has been optimized with a default angle of 45° to the cavity axis, in order to allow a maximal tuning range.
The optimization of the coupler loop, for a cavity loaded with 150 kW of beam leads to an height of 235 mm, which provides a good matching (figure 3.40) and a $S_{1,1}$ computed scattering parameter inferior to -60 dB.
@@ 3.8.9 LLRF Electronics
The LLRF (Low Level RF) controls the RF amplitude and phase of the signal sent to the amplifiers. It also controls the tuning mechanism in order to maintain the resonant frequencies in the cavities. The amplitude and phase control is based on feed-back signals coming from the cavities and amplifiers. In order to avoid multipactoring, the LLRF manages also a pulsing mode that provides for an easy start-up and reduces the power dissipated in the amplifiers. One LLRF system per cavity is planned.
@@
3.8.10 Risks and Mitigation
Risk: Design and Manufacturing Process. The Q factor and eigenfrequency depend on the manufacturing process of the RF Cavity, e.g., surface quality and material purity. Furthermore, there is a risk of multipactoring during run-time as electrons can be accelerated back an forth in the RF field.
Mitigation: A 30% margin was built into the RF power estimates, based on operational experience by IBA. An appropriate tuning range is built into the cavity design. A full scale RF cavity will be built and tested before the start of the construction of the other three cavities. Options of cavity coating exist in case of multipactoring, e.g., Aquadag.
3.9 Cyclotron Particle Simulations
It is necessary to perform an in-depth particle simulation of the primary acceleration of the beam to ensure that the beam reaches energy, maintains quality, and is possible to extract. We do this using OPAL  [adelmann_opal_2019] which is a particle accelerator simulation suite developed at the Paul Scherrer Institute that accounts for space-charge effects. It is based on particle-in-cell (PIC) calculations that track a ’bunch’ of particles through time-varying magnetic and electric fields. The temporal integration within OPAL is carried out using a fourth-order Runge-Kutta method.There is a mode within OPAL, OPAL-cycl, which is designed to simulate cyclotrons. This makes it well-suited for the primary acceleration phase of the IsoDAR cyclotron. Conveniently, OPAL-cycl has built-in objects called PROBEs, that can provide beam diagnostics, and and COLLIMATORs, which can provide halo reduction.
3.9.1 Fieldmaps
A critical aspect of these simulations is that they accurately simulate the electric and magnetic fields within the cyclotron. We performed these simulations using a 1D RF model derived from the voltages provided in Table 3.11. The radially increasing voltage leads to a roughly constant turn separation, which is important for ease of extraction. We conducted these simulations using a 2D median plane magnetic field that is derived from a physically realistic cyclotron model produced by IBA. Of note, this magnetic field map does not include a $\nu_{r}=1$ resonance, which would increase the inter-turn separation in the last few turns at the price of reduced beam quality. This magnetic field is significantly more realistic than our previous work because the cyclotron model allots space for the injection components.
3.9.2 Probes
Probes are a built-in OPAL object that records whenever a particle passes through the region defined by the probe. These particles are simply recorded, and the trajectories of the particles are not affected by their intersection with the probe. This is helpful for determining the “extractability” of the beam. The probe collects information about the position and the momenta of the passing particles. By placing a probe at an azimuth of 25 degrees, the angle at which our electrostatic channel starts, we are able to determine the overlap between sequential turns of the beam path. This is necessary because we want to minimize this overlap for ease of extraction. The overlap between beams provides a lower bound for the power that will be deposited on the electrostatic septum, so we aim to minimize this value. The data from the probe is displayed in Fig 3.42.
3.9.3 Simulation Results
Using the field maps as discussed earlier and the set of 15 collimators as described in the previous sections. We are able to smoothly accelerate the beam up to 60 MeV/amu as shown in Fig 3.43. Additionally, there is minimal growth in beam size in the collimated beam as shown in Fig 3.41. Additionally, a probe at an azimuth of 25 degrees sets a lower limit for extraction at 120W as seen in Fig 3.42.
Figure 3.41: The collimated beam had a much reduced radial RMS spread.
Figure 3.42: There are two potential extraction points
Figure 3.43: There is a smooth acceleration of the beam
3.9.4 Electrostatic Extraction Channels
The first step in our extraction is the construction of a radially directed electric field that peels the last turn away from the second to last turn. This is done by placing a series of thin (0.2 mm) grounded electrodes between the last two turns and a larger puller electrode that is at high negative potential outside of the last turn. This generates a well-shielded electric field in the region between the puller and septum. A rendering of the electrostatic septum is shown in Fig 3.44.There are two such electrostatic channels that increase separation between the last two turns.
Figure 3.44: CAD rendering of electrostatic septum. From Ref. [winklehner_new_2022]
We determine the exact location of each element of the extraction channel iteratively. The first step is to run a Python script on a pre-existing OPAL simulation of the last three turns that determines the initial positions of the septum and the puller. We choose to place the septum evenly between the last two turns. A .txt file is generated that allows for a model of the septum to be built in Autodesk Inventor. With this geometry, we use COMSOL to determine the electric field and add it to the pre-existing RF field of the cyclotron in OPAL. We model the physical dimensions of the septum as COLLIMATOR objects in OPAL. This process is repeated twice at which point further iteration does not result in any improvements in ESC performance. Following this we were able to reduce our losses on the first septum to 205 W which is 2.5% greater than our 200 W goal and we lose 21 W on the second septum. In the future we plan to reintroduce a resonance into the magnetic field which will increase the inter-turn separation at extraction which will greatly reduce this number. The end result of the septum optimization process is shown in Fig 3.45.
Figure 3.45: This is the tracking of 100 particles as they pass through the electrostatic channels (shown in green). It also shows the location of our magnetic extraction channel in grey.
3.9.5 Magnetic Extraction Channels
After the electrostatic extraction channels, the beam will rapidly de-focus as it passes through the radially decreasing magnetic field. In order to avoid this issue, we must include focusing elements that can counteract the effects of the fringe fields of the cyclotron. Due to space constraints, it is not possible to have an active element such as a quadrupole within the cyclotron. Instead, we will use the standard strategy of machining and placing iron bars along the extracted beam path. These bars allow us to locally reshape the magnetic field and are very compact. We model our magnetic extraction off of the MSU cyclotron [MSU_Cyclotron] which generates a strong radially increasing magnetic field that serves to refocus the beam after the ESC. By using the gradients that are produced at MSU as a reference point, we conservatively set a gradient of 1.5 kG/cm which is much less than the 5 kG/cm that they were able to achieve. We directly modified our magnetic field to contain these elements. It is visible in Fig 3.46 that a fairly week magnetic channel is sufficient to maintain beam quality. We do not anticipate difficulties with these channels because the electrostatic channels generate the majority of the turn separation.
Figure 3.46: A 1.5 kG/cm gradient after the electrostatic channel dramatically reduces radial beam spread
3.9.6 Risks and Mitigation
Risk: Inadequate turn separation. If beam bunches are not radially separated by a sufficient amount, unacceptable beam loss will occur on the thin septum rendering its lifetime to be inadequately short.
Mitigation: Understanding beam dynamics sufficiently well that this does not happen, and having adequate beam diagnostics to be able to tune the beam for maximum turn separation. Further, we could place a narrow carbon foil in front of teh septum, dissociating the few $\mathrm{H}_{2}^{+}$ ions that would otherwise hit the septum. The resulting protons have lower magnetic rigidity and bend away from the septum before hitting it. They follow a specific trajectory that can guide them outside the cyclotron to be dumped in a controlled way.
Risk: Placement of the narrow stripping foil. It is not obvious where the optimum location is for the thin stripper. It must be such that stripped protons do not strike the septum, but also that halo particles do not miss the stripper and strike the septum.
Mitigation: Adequate beam dynamics simulations, including tracing of particle orbits.